# Matching with Generalized Sequential Reserves

### v2

@@ 6. Conclusion
This paper introduced a new family of choice rules called Generalized Sequential (GS) choice rules for institutions with multiple divisions. The GS framework unifies and extends various choice rules that appear naturally in real-world market design applications, particularly those involving complex affirmative action policies. We showed that when institutions use GS choice rules, the Cumulative Offer Mechanism (COM) is the unique mechanism that satisfies both stability and strategy-proofness. This uniqueness result provides strong theoretical support for using COM in settings with complex distributional constraints.
We demonstrated that prominent choice rules used in practice, such as aggregate choice rules that use meritorious horizontal choice rules in each vertical category in India and several choice rules used in high school admissions in numerous cities in China, are members of the GS family. We provide a unified theoretical framework for analyzing them.
The framework developed in this paper provides a foundation for analyzing complex matching markets with distributional constraints. By unifying various existing approaches under the GS umbrella, it opens new avenues for both theoretical analysis and practical market design.
@@ 7. Appendix
Since $Y_{j}=Z_{j}$ for all $j=1,\ldots,k-1$, we obtain $r_{j}=\overline{r}_{j}$ for all $j=1,\ldots,k-1$. Then $q^{k}_{s}(r_{1},...,r_{k-1})=\overline{q}_{s}^{k}(\overline{r}_{1},...,\overline{r}_{k-1}).$ Since there are no two contracts of an individual chosen by $\overline{C}_{s}$, all the remaining contracts of individuals—whose contracts were chosen by previous divisions—are rejected by $C^{k}_{s}(\overline{Y}^{k};\overline{q}_{s}^{k}(\overline{r}_{1},...,\overline{r}_{k-1}))$. Therefore, the IRC of $C^{k}_{s}$ implies that $C^{k}_{s}(\overline{Y}^{k};\overline{q}_{k}^{s}(\overline{r}_{1},...,\overline{r}_{k-1}))=C_{s}^{k}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1})).$ Thus, we obtain $\overline{Y}_{k}=Y_{k}$, $\overline{r}_{k}=r_{k}$, and $\overline{q}^{k+1}_{s}(\overline{r}_{1},...,\overline{r}_{k})=q^{k+1}_{s}(r_{1},...,r_{k})$. Since each division selects the same set of contracts under $C_{s}$ and $\overline{C}_{s}$, the result follows.
@@
Since the choice function of division 1 satisfies the IRC, we have $Y_{1}=Z_{1}$. Moreover, $r_{1}=\tilde{r}_{1}$ and $Y^{2}\setminus\{x\}=Z^{2}$. By induction, for each $j=2,...,k-1$, assume that
$Y_{j}=Z_{j},\;r_{j}=\tilde{r}_{j},\;and\;Y^{j}\setminus\{x\}=Z^{j}.$
@@
By our inductive assumption that $r_{j}=\tilde{r}_{j}$ for each $j=2,...,k-1$, division $k$’s capacity is the same under both choice processes. The number of remaining slots is the same as well. That is, $r_{k}=\tilde{r}_{k}$.
Finally, we know that $x$ is not chosen from the set $Z^{k}\cup\{x\}$, then we have $Y^{k+1}=Z^{k+1}\cup\{x\}.$
Since for all $j\in\{1,...,K_{s}\}$, $Y_{j}=Z_{j}$, we have $\overline{C}_{s}(Y;q_{s})=\overline{C}_{s}(Z;q_{s})$. Thus, $\overline{C}_{s}$ satisfies the IRC.
@@
Consider $Y\subseteq X$ such that $Y\neq\overline{C}_{s}(Y;q_{s})$. Let $x$ be such that $x\in Y\setminus\overline{C}_{s}(Y;q_{s})$, and let $z$ be an arbitrary contract in $X\setminus Y$. Let $Z=Y\cup\{z\}$. We need to show that $x\notin\overline{C}_{s}(Z;q_{s}).$
Let $(Y_{j},r_{j},Y^{j})$ and $(Z_{j},\tilde{r}_{j},Z^{j})$ be sequences of the set chosen contracts, the number of vacant slots, and the set of remaining contracts of the divisions $j=1,...,K_{s}$ from $Y$ and $Z$, respectively, under $\overline{C}$.
If $z\in Z\setminus\overline{C}_{s}(Z;q_{s})$, then the IRC of $\overline{C}_{s}$ implies $\overline{C}_{s}(Z;q_{s})=\overline{C}_{s}(Y;q_{s})$. Therefore, $x\notin\overline{C}_{s}(Z;q_{s})$.
Suppose $z\in\overline{C}_{s}(Z;q_{s})$. Consider the first division. If $z\notin Z_{1}$, then by IRC we have $Y_{1}=Z_{1}$ and, thus, $r_{1}=\widetilde{r}_{1}$. Moreover, we obtain $Z^{2}=Y^{2}\cup\{z\}$ and $q^{2}_{s}(r_{1})=q^{2}_{s}(\widetilde{r}_{1})$. Now, consider the case where $z\in Z_{1}$. Note that, by the substitutes condition, all contracts that are rejected from $Y^{1}$ are rejected from $Z^{1}$. By the size monotonicity condition, $\mid Z_{1}\mid\geq\mid Y_{1}\mid$. There are two possibilities:
@@
    $Z_{1}=Y_{1}\cup\{z\}\setminus\{w\}$ for some $w\in Y_{1}$.
In the former case, we obtain $Y^{2}=Z^{2}$ and $r_{1}=1+\widetilde{r}_{1}$. Moreover, we have $q^{2}_{s}(\widetilde{r}_{1})\leq q^{2}_{s}(r_{1})\leq 1+q^{2}_{s}(\widetilde{r}_{1})$ due to the monotonicity of capacity transfers. In the latter case, we obtain $Z^{2}=Y^{2}\cup\{w\}$ and $r_{1}=\widetilde{r}_{1}$. Therefore, we have $q^{2}_{s}(r_{1})=q^{2}_{s}(\widetilde{r}_{1})$.
Suppose now that for all $\gamma=j,...,k-1$ we have that either
$\left[Z^{\gamma}=Y^{\gamma}\cup\{w\}\;for\;some\;w\;and\;q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})=q^{\gamma+1}_{s}(r_{1},...,r_{\gamma})\right]$ (1)
@@
$\left[Z^{\gamma}=Y^{\gamma}\;and\;q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})\leq q^{\gamma+1}_{s}(r_{1},...,r_{\gamma})\leq 1+q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})\right].$ (2)
We have already shown that it holds for $\gamma=1$ and next we show that it also holds for $\gamma=k$.
@@
If $w$ is chosen from $Z^{k-1}$, we have two possibilities, depending on whether the capacity of division $k$ is exhausted in the choice process starting with $Y^{0}$ or not. If it is not exhausted, then we obtain
@@
The monotonicity of the transfer function implies
@@
Note that the first inequality follows from the fact that $\tilde{r}_{i}\leq r_{i}$ for all $i=1,...,k$. The second inequality follows from the second condition of the monotonicity of the transfer function.
[…]
We now analyze the latter case, in which we have $Z^{k-1}=Y^{k-1}$ and either $q^{k}_{s}(r_{1},...,r_{k-1})=q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$ or $q^{k}_{s}(r_{1},...,r_{k-1})=1+q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$. If $q_{s}^{k}(r_{1},...,r_{k-1})=q_{s}^{k}(\tilde{r}_{1},...,\tilde{r}_{k-1})$, then, given that $Z^{k-1}=Y^{k-1}$, we have $Z^{k}=Y^{k}$. This also implies $r_{k}=\tilde{r}_{k}$. Moreover, we obtain $q^{k+1}_{s}(r_{1},...,r_{k})=q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})$ by the monotonicity of the transfer function. Note that $\tilde{r}_{i}\leq r_{i}$ for all $i=1,...,k$ implies $q^{k+1}_{s}(r_{1},...,r_{k})\geq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})$ by the first monotonicity condition. The second condition of monotonicity implies $q^{k+1}_{s}(r_{1},...,r_{k})\leq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})$.
@@
Case 1: If $C_{s}^{k}(Z^{k-1};q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))=C_{s}^{k}(Y^{k-1};q^{k}_{s}(r_{1},...,r_{k-1})),$ then we have $Z^{k}=Y^{k}$. In addition, the monotonicity of transfer function implies
@@
Since $x\notin Y_{k}$, we have $x\notin Z_{k}$ for all $k=1,...,K_{s}$. We conclude that $x\notin\overline{C}_{s}(Y\cup\{z\};q_{s})$.
@@
Suppose that $\tilde{r}_{j}\leq r_{j}$ and $Y^{j}\subseteq Z^{j}$ hold for all $j=1,...,k-1$. We need to show that
@@
We first show (1). Given that $\tilde{r}_{j}\leq r_{j}$ for all $j=1,...,k-1$, the first condition of monotonicity implies that $q^{k}_{s}(r_{1},...,r_{k-1})\geq q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$. By the size monotonicity, we have
@@
That is $\tilde{r}_{k}\leq r_{k}$.
We now prove (2). Consider a contract $x\in Y^{k}$. That is, $x\notin C_{s}^{k}(Y^{k-1},q^{k}_{s}(r_{1},...,r_{k-1}))$. When the set $Y^{k-1}$ is expanded to $\tilde{Y}^{k-1}$, $x$ cannot be chosen since the choice function satisfies the substitutes condition. That is,
@@
Now, let $\tau_{j}=r_{j}-\tilde{r}_{j}$. As shown above, $\tau_{j}\geq 0$ for all $j=1,...,K_{s}$. Putting in $r_{j}=q_{s}^{j}(r_{1},...,r_{j-1})-\mid Y_{j}\mid$ and $\tilde{r}_{j}=q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{k-1})-\mid Z_{j}\mid$ gives
@@
Since $q_{s}^{j}(r_{1},...,r_{j-1})\geq q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{j-1})$ for all $j=2,...,K_{s}$, we obtain
@@
That is, $\mid\overline{C}_{s}(Y;q_{s})\mid\leq\mid\overline{C}_{s}(Z;q_{s})\mid$, the desired conclusion.
@@
We need to show that $C_{s}^{mT}$ is in the GS family. The transfer functions under $C_{s}^{mT}$ satisfy monotonicity in a straightforward way. Fleiner (2001) shows that $C^{m}$ satisfies the substitute property Sönmez and Yenmez (2022) prove that $C^{m}$ satisfies size monotonicty.
We then only need to show that $C^{m}$ satisfies quota-monotonicity. That is, for any given set of applicants $A\subseteq\mathcal{I}$ and $q\in\mathbb{N}$, we need to show that
@@
Without loss of generality, let $h_{j}$ be the horizontal type whose capacity increases by 1. For each $i\in C^{m}\left(A,q\right)$, given the applicants chosen before $i$ in the first step of $C^{m}$, $i$ is chosen because it increases the utilization of horizontally reserved positions, which means that horizontally reserved positions are available for her. Adding one more type $h_{j}$ position (weakly) increases the horizontally reserved positions available to her. Thus, $i$ continues to be chosen. That is, $C^{m}\left(A,q\right)\subseteq C^{m}\left(A,q+1\right)$. Since each applicant chosen when capacity was $q$ continues to be chosen when capacity increases by one, the number of applicants chosen in Step 1 of $C^{m}$ stays the same or increases by one. By Step 2 of $C^{m}$, we can conclude that $\mid C^{m}\left(A,q+1\right)\mid-\mid C^{m}\left(A,q\right)\mid\leq 1$.

### v3

@@ 6. Conclusion
This paper introduced a new family of choice rules for institutions with multiple divisions. The GSq framework unifies and extends various choice rules that appear naturally in real-world market design applications, particularly those involving complex affirmative action policies. We showed that when institutions use GSq choice rules, the Cumulative Offer Mechanism (COM) is the unique mechanism that satisfies both stability and strategy-proofness. This uniqueness result provides strong theoretical support for using COM in settings with complex distributional constraints.
We demonstrated that prominent choice rules used in practice, such as aggregate choice rules that use meritorious horizontal choice rules in each vertical category in India and several choice rules used in high school admissions in numerous cities in China, are members of the GSq family. We provide a unified theoretical framework for analyzing them.
The framework developed in this paper provides a foundation for analyzing complex matching markets with distributional constraints. By unifying various existing approaches under the GSq family, it opens new avenues for both theoretical analysis and practical market design.
@@ 7. Appendix
Since $Y_{j}=Z_{j}$ for all $j=1,\ldots,k-1$, we obtain $r_{j}=\overline{r}_{j}$ for all $j=1,\ldots,k-1$, which implies by the definition of the transfer function
$q^{k}_{s}(r_{1},...,r_{k-1})=\overline{q}_{s}^{k}(\overline{r}_{1},...,\overline{r}_{k-1}).$
Since there are no two contracts of an individual chosen by $\overline{C}_{s}$, all the remaining contracts of individuals—whose contracts were chosen by previous divisions—are rejected by $C^{k}_{s}(\overline{Y}^{k};\overline{q}_{s}^{k}(\overline{r}_{1},...,\overline{r}_{k-1}))$. Therefore, the IRC of $C^{k}_{s}$ implies that [^17] Divisions’ choice rules are assumed to satisfy the substitutes property and size monotonicity. The conjugation of the two implies the irrelevance of rejected contracts (IRC) condition.
$C^{k}_{s}(\overline{Y}^{k};\overline{q}_{k}^{s}(\overline{r}_{1},...,\overline{r}_{k-1}))=C_{s}^{k}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1})).$
Thus, we obtain $\overline{Y}_{k}=Y_{k}$, $\overline{r}_{k}=r_{k}$, and $\overline{q}^{k+1}_{s}(\overline{r}_{1},...,\overline{r}_{k})=q^{k+1}_{s}(r_{1},...,r_{k})$. Since each division selects the same set of contracts under $C_{s}$ and $\overline{C}_{s}$, the result follows.
In the following, we prove that the rule $\overline{C}_{s}$ satisfies the irrelevance of rejected contracts (IRC) condition, substitutes property, and size monotonicity.
@@
Since the choice function of division 1 satisfies the IRC, we have $Y_{1}=Z_{1}$. Moreover, $r_{1}=\tilde{r}_{1}$ and $Y^{2}\setminus\{x\}=Z^{2}$. By induction, for each j=2,,k1j=2,...,k-1, assume that
$Y_{j}=Z_{j},\;r_{j}=\tilde{r}_{j},\;\text{and}\;Y^{j}\setminus\{x\}=Z^{j}.$
@@
By our inductive assumption that $r_{j}=\tilde{r}_{j}$ for each j=2,,k1j=2,...,k-1, division $k$’s capacity is the same under both choice processes. The number of remaining slots is the same as well. That is, $r_{k}=\tilde{r}_{k}$.
Moreover, we know that $x$ is not chosen from the set $Z^{k}\cup\{x\}$, then we have
$Y^{k+1}=Z^{k+1}\cup\{x\}.$
Since for all $j\in\{1,...,K_{s}\}$, $Y_{j}=Z_{j}$, we conclude $\overline{C}_{s}(Y;q_{s})=\overline{C}_{s}(Z;q_{s})$.
@@
Consider $Y=Y^{0}\subseteq X$ such that $Y\neq\overline{C}_{s}(Y;q_{s})$. Let $x$ be such that $x\in Y\setminus\overline{C}_{s}(Y;q_{s})$, and let $z$ be an arbitrary contract in $X\setminus Y$. Let $Z=Z^{0}=Y\cup\{z\}$. We need to show that $x\notin\overline{C}_{s}(Z;q_{s}).$
Let $(Y_{j},r_{j},Y^{j})$ and $(Z_{j},\tilde{r}_{j},Z^{j})$ be sequences of the set chosen contracts, the number of vacant slots, and the set of remaining contracts of the divisions $j=1,...,K_{s}$ from $Y$ and $Z$, respectively, under $\overline{C}_{s}$.
If $z\notin\overline{C}_{s}(Z;q_{s})$, then the IRC of $\overline{C}_{s}$ implies $\overline{C}_{s}(Z;q_{s})=\overline{C}_{s}(Y;q_{s})$. Thus, $x\notin\overline{C}_{s}(Z;q_{s})$.
Now suppose $z\in\overline{C}_{s}(Z;q_{s})$. Let $k^{*}$ be the division in which $z$ is chosen.
Since $z\notin Z_{j}$, by the IRC of the choice rules, we have $Z_{j}=Y_{j}$ for $j=1,...,k^{*}-1$. Thus, we obtain $r_{j}=\tilde{r}_{j}$, for $j=1,...,k^{*}-1$. Moreover, for each $j=1,...,k^{*}-1$, the division capacities are the same under the two choice processes, starting with $Y$ and $Z$, respectively. That is,
$q_{s}^{j}(r_{1},...,r_{j-1})=q_{s}^{j}(\widetilde{r}_{1},...,\widetilde{r}_{j-1}).$
Now, we consider division $k^{*}$. We know $z\in Z_{k^{*}}=C_{s}^{k^{*}}(Z^{k^{*}-1};q_{s}^{k^{*}}(\widetilde{r}_{1},...,\widetilde{r}_{k^{*}-1}))$.
By the substitutes property of $C_{s}^{k^{*}}$, all contracts that are rejected from $Y^{k^{*}-1}$ are rejected from $Z^{k^{*}-1}=Y^{k^{*}-1}\cup\{z\}$. By size monotonicity, $\mid Z_{k^{*}}\mid\geq\mid Y_{k^{*}}\mid$. There are two possibilities:
@@
    $Z_{k^{*}}=Y_{k^{*}}\cup\{z\}\setminus\{w\}\text{for some }w\in Y_{k^{*}}$.
In the former case, we obtain $Y^{k^{*}}=Z^{k^{*}}$ and $r_{k^{*}}=1+\widetilde{r}_{k^{*}}$. Moreover, by the monotonicity of capacity transfers, we have
$q^{k^{*}+1}_{s}(\widetilde{r}_{1},...,\widetilde{r}_{k^{*}})\leq q^{2}_{s}(r_{1},...,r_{k^{*}})\leq 1+q^{2}_{s}(\widetilde{r}_{1},...,\widetilde{r}_{k^{*}}).$
In the latter case, we obtain $Z^{k^{*}}=Y^{k^{*}}\cup\{w\}$ and $r_{k^{*}}=\widetilde{r}_{k^{*}}$. Since $r_{j}=\widetilde{r}_{j}$ for $j=1,...,k^{*}$, we thus have
$q^{k^{*}+1}_{s}(r_{1},...,r_{k^{*}})=q^{k^{*}+1}_{s}(\widetilde{r}_{1},...,\widetilde{r}_{k*}).$
Suppose now that for all $\gamma=k^{*},...,k-1$, one of the following hold.
$\left[Z^{\gamma}=Y^{\gamma}\cup\{w\}\;\text{for some}\;w\;\text{and}\;q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})=q^{\gamma+1}_{s}(r_{1},...,r_{\gamma})\right]$ (5)
@@
$\left[Z^{\gamma}=Y^{\gamma}\;\text{and}\;q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})\leq q^{\gamma+1}_{s}(r_{1},...,r_{\gamma})\leq 1+q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})\right].$ (6)
We have already shown that it holds for $\gamma=k^{*}$. Assuming that one of them holds for $\gamma=k^{*},...,k-1$, we need to show that one of them holds for $\gamma=k$.
@@
If $w\in Z_{k}$, we have two possibilities, depending on whether the capacity of division $k$ is exhausted in the choice process starting with $Y^{0}$ or not. If it is not exhausted (that is, $|Y_{k}|<q_{s}^{k}(r_{1},...,r_{k-1})$), then we obtain
@@
The monotonicity of the transfer function implies the following.
@@
Note that the first inequality follows from the first condition of the monotonicity of transfer function and the observation that $\tilde{r}_{i}\leq r_{i}$ for all i=1,,ki=1,...,k. The second inequality follows from the second condition of the monotonicity of the transfer function. Note that $|Z_{1}|+\cdots+|Z_{k}|=1+|Y_{1}|+\cdots+|Y_{k}|$.
If the capacity of the division $k$ is exhausted (that is, $|Y_{k}|=q_{s}^{k}(r_{1},...,r_{k-1})$), then $w\in Z_{k}$ implies that there exists a contract $\upsilon$ that is chosen from $Y^{k-1}$ but rejected from $Z^{k-1}$. Then, we obtain $Z^{k}=Y^{k}\cup\{\upsilon\}$ since $C_{s}^{k}$ satisfies the substitutes property and IRC.
In addition, the capacity of division $k$ is the same in both processes. In this case, we have $r_{k}=\tilde{r}_{k}=0$. Since $\tilde{r}_{i}=r_{i}$ for all i=1,,ki=1,...,k, we obtain, by the definition of the transfer policy, $q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})=q^{k+1}_{s}(r_{1},...,r_{k})$.
We now analyze the latter case, in which condition (6) holds. That is, we have $Z^{k-1}=Y^{k-1}$ and either $q^{k}_{s}(r_{1},...,r_{k-1})=q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$ or $q^{k}_{s}(r_{1},...,r_{k-1})=1+q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$.
If $q_{s}^{k}(r_{1},...,r_{k-1})=q_{s}^{k}(\tilde{r}_{1},...,\tilde{r}_{k-1})$, then, given that $Z^{k-1}=Y^{k-1}$, we have $Z^{k}=Y^{k}$. This also implies $r_{k}=\tilde{r}_{k}$. Note that given $\tilde{r}_{i}\leq r_{i}$ for all i=1,,ki=1,...,k, the first condition of the monotone transfer function implies
$q^{k+1}_{s}(r_{1},...,r_{k})\geq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k}).$ (7)
Moreover, from the second condition of monotonicity, we can write
$\sum_{m=2}^{k+1}[q_{s}^{m}(r_{1},...,r_{m-1})-q_{s}^{m}(\widetilde{r}_{1},...,\widetilde{r}_{m-1})]\leq\sum_{m=1}^{k-1}[r_{m}-\widetilde{r}_{m}].$
We can rewrite the above inequality as follows.
$q^{k+1}_{s}(r_{1},...,r_{k})+\sum_{m=1}^{k}[q_{s}^{m}(r_{1},...,r_{m-1})-q_{s}^{m}(\widetilde{r}_{1},...,\widetilde{r}_{m-1})]-\sum_{m=1}^{k-1}[r_{m}-\widetilde{r}_{m}]\leq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k}).$
Rearranging the terms gives us the following.
$q^{k+1}_{s}(r_{1},...,r_{k})+\sum_{m=1}^{k}[q_{s}^{m}(r_{1},...,r_{m-1})-r_{m}]-\sum_{m=1}^{k}[q_{s}^{m}(\widetilde{r}_{1},...,\widetilde{r}_{m-1})-\widetilde{r}_{m}]\leq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k}).$
The terms in the summations are the number of contracts in the chosen sets $Y_{m}$ and $Z_{m}$, respectively. Plugging them in gives us the following.
$q^{k+1}_{s}(r_{1},...,r_{k})+\sum_{m=1}^{k}|Y_{m}|-\sum_{m=1}^{k}|Z_{m}|\leq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k}).$
Note that $\sum_{m=1}^{k}|Y_{m}|-\sum_{m=1}^{k}|Z_{m}|=-1$. Because we have $|Z|=|Y|+1$ and $Z^{k}=Y^{k}$.
Thus, we obtain
$q^{k+1}_{s}(r_{1},...,r_{k})\leq 1+q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k}).$ (8)
Inequalities (7) and (8) together imply
@@
Case 1: If $C_{s}^{k}(Z^{k-1};q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))=C_{s}^{k}(Y^{k-1};q^{k}_{s}(r_{1},...,r_{k-1})),$ then we have $Z^{k}=Y^{k}$. In addition, the monotonicity of the transfer function implies
@@
To conclude the proof, since either (5) or (6) holds in each division, given that $x\notin Y_{k}$, we have $x\notin Z_{k}$ for all $k=1,...,K_{s}$. We conclude that $x\notin\overline{C}_{s}(Y\cup\{z\};q_{s})$. Thus, $\overline{C}_{s}$ satisfies the substitutes property.
@@
Suppose that $\tilde{r}_{j}\leq r_{j}$ and $Y^{j}\subseteq Z^{j}$ hold for all j=1,,k1j=1,...,k-1. We need to show that
@@
We first show (1). Given that $\tilde{r}_{j}\leq r_{j}$ for all j=1,,k1j=1,...,k-1, the first condition of monotonicity implies that $q^{k}_{s}(r_{1},...,r_{k-1})\geq q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$. By the size monotonicity, we have
@@
That is,
$\tilde{r}_{k}\leq r_{k}.$
We now prove (2). Consider a contract $x\in Y^{k}$. That is, $x\notin C_{s}^{k}(Y^{k-1},q^{k}_{s}(r_{1},...,r_{k-1}))$. When the set $Y^{k-1}$ is expanded to $\tilde{Y}^{k-1}$, $x$ cannot be chosen since the choice function satisfies the substitutes property. That is,
@@
Now, let $\tau_{j}=r_{j}-\tilde{r}_{j}$. As shown above, $\tau_{j}\geq 0$ for all $j=1,...,K_{s}$. Putting in
$r_{j}=q_{s}^{j}(r_{1},...,r_{j-1})-\mid Y_{j}\mid$
and
$\tilde{r}_{j}=q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{k-1})-\mid Z_{j}\mid$
we obtain the following.
@@
Since $q_{s}^{j}(r_{1},...,r_{j-1})\geq q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{j-1})$ for all $j=2,...,K_{s}$, we obtain the following.
@@
That is,
$\mid\overline{C}_{s}(Y;q_{s})\mid\leq\mid\overline{C}_{s}(Z;q_{s})\mid.$
@@
We need to show that $C_{s}^{mT}$ is in the GSq family. The transfer functions under $C_{s}^{mT}$ satisfy monotonicity because $q_{s}^{D}=r_{OBC}$. fleiner2001matroid shows that the first Step 1 of $C^{m}$ satisfies the substitute property. Moreover, sonmez2022affirmative prove that $C^{m}$ satisfies size monotonicity. We then only need to show that $C^{m}$ satisfies quota-monotonicity. That is, for any given set of applicants $A\subseteq\mathcal{I}$ and $q\in\mathbb{N}$, we need to show that
@@
Let $h_{j}$ be the horizontal type whose capacity increases by 1, w.l.o.g. Consider Step 1 (the Greedy process) under $C^{m}(A,q)$ and $C^{m}(A,q+1)$. Let $n^{*}$ be the first step in which the two choice procedures diverge. That is, the $n^{*}$-highest scoring individual is not selected under $C^{m}(A,q)$, but is chosen under $C^{m}(A,q+1)$. Note that in Steps $1.1$ to $1.(n^{*}-1)$, the same set of students are chosen in $C^{m}(A,q)$ and $C^{m}(A,q+1)$. If $i_{n^{*}}\notin C^{m}(A,q)$ and $i_{n^{*}}\in C^{m}(A,q+1)$, then it must be the case that $i_{n^{*}}$ is matched to a position reserved for the horizontal type $h_{j}$.
Note also that the set of remaining students is and the set of horizontally reserved positions are the same at the end of Step $1.n^{*}$ of the greedy algorithm. Thus, the in the remaining steps, the same set of students will be chosen. Thus, each student who is chosen under $C^{m}(A,q)$ will continues to be chosen under $C^{m}(A,q+1)$. That is,
$C^{m}\left(A,q\right)\subseteq C^{m}\left(A,q+1\right).$
Since each applicant chosen when capacity was $q$ continues to be chosen when capacity increases by one, the number of applicants chosen in Step 1 of $C^{m}$ stays the same or increases by one. By Step 2 of $C^{m}$, we can conclude that
Proof of Proposition 2
The choice rules used in Step 1 and Step 2 are both q-responsive induced by applicants test scores. Thus, they satisfy the substitutes property, size monotonicity, and quota monotonicity. The capacity of the open category positions that are allocated in Step 2 is $q_{s}^{o}+\sum_{m\in M}r^{m}_{s}$ where $r_{s}^{m}$ denotes the number of unfilled seats that were reserved for middle school $m\in M$. This transfer policy straightforwardly satisfy monotonicity condition. Thus, $\mathcal{C}^{SimRO}$ is a GSq choice rule according to Definition 3.
Proof of Proposition 3
The choice rules used in Steps 1-3 are all q-responsive induced by test scores. Therefore, they satisfy the substitutes property, size monotonicity, and quota monotonicity. The capacity of the open category positions that are allocated in Step 3 is $\sum_{m\in M}r^{m}_{s}$, where $r_{s}^{m}$ denotes the number of unfilled seats that were reserved for middle school $m\in M$. This transfer policy is monotonic. Then, by Definition 3, $\mathcal{C}^{SimORO}$ is a GSq choice rule.
Proof of Proposition 4
Given $I,J\subset\mathcal{I}$ such that $I\cap J=\emptyset$, $\mathcal{C}^{SimSep}(I,J)$ assigns the top $q_{s}^{o}$ applicants within the set $I$ to open positions and the top $q_{s}^{m}$ applicants within $J\cap\mathcal{I}^{m}$ to reserved positions for $m\in M$.
$\mathcal{C}^{SimSep}$ satisfies the subsitutes condition and size monotonicity. Thus, it belongs to the GSq family.
Proof of Proposition 5
Given $I,J\subset\mathcal{I}$ such that $I\cap J=\emptyset$, $\mathcal{C}^{SimFlex}(I,J)$ assigns the top $q_{s}^{m}$ applicants within the set $J$ to positions reserved for $m\in M$ in the first step. In the second step, it considers all remaining positions and assigns the top $q_{s}^{o}$ applicants within $I$ to open positions.
$\mathcal{C}^{SimSep}$ satisfies the subsitutes condition and is q-acceptant. Hence, it satisfies size monotonicity. Therefore, it belongs to the GSq family.
Proof of Proposition 6
hu2025verifiable show that $\mathcal{C}^{SimOR}$ satisfies the subsititues condition and is $q$-acceptant. Since $q$-acceptance implies size monotonicity, it also satisfies size monotonicity. Since $\mathcal{C}^{SimOR}$ satisfies the substitutes property and size monotonicity, it can be regarded as a single division’s choice rule. Thus, it is a GSq choice rule.